OpenStudy (anonymous):

A pilot can travel 500 miles with the wind in the same amount of time as 340 miles against the wind. Find the speed of the wind if the pilots speed in still air is 210 mph. please help!

OpenStudy (anonymous):

@AlexandervonHumboldt2 @dan815 @Joel_the_boss @Val14

OpenStudy (anonymous):

@Jamielynne @josedavid @Elwell98 @Quan99 @k12andstudyislandhelp

OpenStudy (anonymous):

idk srry

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

@bohotness @fashionlove @Bee_see @blah124 @KierseyClemons @bibby

OpenStudy (anonymous):

@rachelledelija123 @Squirrels @sissyedgar @W.W.G.D @esam2 @linn99123 @Quan99

OpenStudy (anonymous):

@NaomiBell1997 @nothingwasthesame @goatgal102

OpenStudy (anonymous):

@InExileWeTrust @yoyogators @owls2001 @xokatexo @zach1751

OpenStudy (anonymous):

@Answers101

OpenStudy (sissyedgar):

I tink its 370! Bbt im not sure!

OpenStudy (sissyedgar):

but*

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

nope :( sadly it's not :(

OpenStudy (anonymous):

i have 2 more trys tho

OpenStudy (anonymous):

do u know what else it might be?

OpenStudy (anonymous):

@Gebooors @CloverRacer @CausticSyndicalist @mathmath333 @luckycoins888 @MikeyMaximum

OpenStudy (luckycoins888):

lemme think for a min k?

OpenStudy (luckycoins888):

try 120

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no, now i have one more try... @luckycoins888

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

@mwill<3 @annon1123 @Molly_Qian @Loser66 @sleepyjess

OpenStudy (anonymous):

huh?

OpenStudy (bohotness):

:d yes you rright :D

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \normalsize \text{let the speed of the wind be } x\quad mph\hspace{.33em}\\~\\ \normalsize \text{let the speed of the plane in still air be } y\quad mph\hspace{.33em}\\~\\ \normalsize \text{pilot speed with the wind}=(x+y) \quad mph\hspace{.33em}\\~\\ \normalsize \text{pilot speed against the wind}=(y-x) \quad mph\hspace{.33em}\\~\\ t_{1}=t_2\hspace{.33em}\\~\\ \dfrac{500}{x+y}=\dfrac{340}{y-x}\hspace{.33em}\\~\\ \normalsize \text{but the speed of the plane in still air is given } y=210 \quad mph\hspace{.33em}\\~\\ \dfrac{500}{x+210}=\dfrac{340}{210-x}\hspace{.33em}\\~\\ \normalsize \text{find x} \hspace{.33em}\\~\\ \end{align}}\)